OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (anonymous):

f(x) = x4 - 8x3 - 12x2 + 400x - 1599 f(x) = x4 - 200x2 + 800x - 1599 f(x) = x4 - 98x2 + 800x - 1599 f(x) = x4 - 8x3 + 12x2 - 400x + 1599